A bullet is fired vertically upward (don't) with a muzzle velocity of 671.ft/s.

a) when does if reach max altitude?

b) what is that max altitude (above the gun)?

c) when does it return to hit the gun (or stupid sportsman)?


1
Expert's answer
2022-06-25T09:01:42-0400

a) when does if reach max altitude?

"t_1=\frac{v_0}{g}=\rm\frac{671\: ft/s}{32.17\: ft/s^2}=21\: s"

b) what is that max altitude (above the gun)?

"h_{\max}=\frac{v_0^2}{2g}=\rm\frac{(671\: ft/s)^2}{2*32.17\: ft/s^2}=7000\: ft"

c) when does it return to hit the gun (or stupid sportsman)?

"t_2=2t_1=2*21\: \rm s=42\: s"